OpenStudy (anonymous):

1. The length of a diagonal of a square is 7.5 mm. Find the area of the square. 2. . The length of a diagonal of a square is 5 inches. Find the perimeter of the square

OpenStudy (ajprincess):

 Using pythagoras theorem a^2 can be found a^2+a^2=d^2 2a^2=d^2 a^2=d^2/2 here a-length of a side of the square d-length of the diagonal a^2-area